# On Uncertainty Calibration for Equivariant Functions

### v2

@@ 4.3 Improved Bounds for bi-Lipschitz Invariant Functions
Notice that the upper bounds in Proposition˜1 and Theorem˜4 are in terms of $r(p)$. The density $r(p)$ is not in general easily derivable from $q(x)$. In order to express each bound in terms of $q(x)$, we introduce extra assumptions on $h_{P}$.
From Hörmander (2015), we have that if $h_{P}(x)$ is continuously differentiable and has gradient nowhere 0, then
@@
Assume $h_{P}(x)$ is differentiable, has gradient nowhere 0, and is $(K_{1},K_{2})$ bi-Lipschitz continuous. Let $Gx^{\diamond}$ be the orbit with the least integrated probability density, i.e. $x^{\diamond}=\underset{x\in X}{\arg\min}\int_{Gx}q(x)dx$. Then
@@
We start with the bound from Theorem˜4. Substituting the expression for $r(p)$ from Equation˜5 into the upper bound from Theorem˜4 and using the Lipschitz constant to bound the gradient gives
@@
As in Corollary˜2, the $\int_{P_{2}}dp$ term is unity in the case of binary classification. This enables us to further characterise the utility of the bound in the examples that follow.
@@ 4.4 Invariant Upper Bounds Examples
A natural question is how much tighter is the upper bound on ECE under the assumption of invariance (Theorem˜4) than the bound in the unconstrained case (Proposition˜1). In the case $h_{P}(x)$ is bi-Lipschitz with Lipschitz constants $K_{1}$ and $K_{2}$, Proposition˜2 gives a relatively concrete answer; the gap is $k(Gx^{*})K_{2}\int_{P_{2}}dp\int_{Gx^{\diamond}}q(x)dx$. In order to understand the gap when the function is not assumed to be bi-Lipshitz, we consider several examples with specific $r(p)$. The examples illuminate how the tendency of a model to be uncertain can tighten the bound for functions both with and without invariance. Specifically, Example˜1 considers a binary classification task where $r(p)$ is a truncated normal distribution. We consider means $\mu$ that correspond to low, medium, and high confidence.
@@
Let us now consider the unconstrained bound for three different example distributions $r(p)$, each corresponding to correspond to low, medium, and high confidence.
@@
The following example illustrates the how to use the bound in Corollary˜2 and how tight it is. That is, we are in the special case when we know the which elements $x$ are in a given fiber $\mathcal{F}_{p}$ and the task is binary classification.
@@
Having only two confidence values may be reasonable in real world settings if there is a prevailing noise (e.g., shadow, camera artifacts, background patterns) on just one side of the field of view of a camera, leading to approximately two different confidence regions in our model output.
@@
Consider the same dataset as in Example˜2, but assume $h$ has rotation invariance to SO(2) instead of reflection invariance. In this case, there is only one orbit and so $h_{Y}$ and $h_{P}$ each take one value. There is only one fiber $F_{p}=X$, which is the entire dataset, as illustrated in Figure˜2c. The classification error is minimized when $h$ predicts each label as blue, since there are more blue labels then orange labels in our dataset. This gives an error lower bound of $0.3$, and an ECE upper bound of $0.7$. This is the same result for reflection invariance in the special case where the confidences are the same across each fiber.
Figure 2: Dataset with pointwise incorrect invariance. The first two panels indicate that the model is invariant $C_{2}$ reflections over the $x$-axis. The presence of diagonal lines in the first panel depicts that the model has $2$ confidence fibers. The last panel indicates that the model has rotation invariance. Colors represent labels in all panels.
@@ 4.6 Invariant Lower Bound Example
For Lipschitz functions $f$ and $g$ with Lipschitz constants $L_{1}$ and $L_{2}$, the composition $f(g(x))$ has Lipschitz constant $L_{2}L_{1}$ and the sum $f(x)+g(x)$ has Lipschitz constant $L_{1}+L_{2}$. Moreover, the Lipschitz constant for a linear map is given by its maximum singular value $\sigma_{\rm max}$. Finally, note that ReLU is 11-Lipschitz. Thus, the Lipschitz constant for Equation˜6 is $\sigma_{\rm max}(1^{T})(\sigma_{\rm max}(I)+\sigma_{\rm max}(11^{T}))=\sigma_{\rm max}(1^{T})(1+\sigma_{\rm max}(11^{T}))$. For $n=24$, we can compute these values to be $24$ and approximately $4.9$ respectively.
@@
Thus far, we have derived a lower bound tjat does not depend on the size of the fibers and has no dependence on $|\nabla h_{P}|$. What is left is to calculate the integrated density on $Gx^{*}$ and the accuracy lower bound $m^{\prime}$.
Since there is only one orbit, $Gx^{*}=X$ and the integrated density over $Gx^{*}$ is unity. Moreover, $m^{\prime}$ reduces to the global accuracy lower bound. Noting that $A_{1}=\vec{a}$ will occur with probability $0.25$ gives us our accuracy lower bound $m^{\prime}$. So, we compute
@@ Results.
Figure˜4 shows that ECE does not follow a roughly monotonic improvement as we increase the cyclic group order in the same way accuracy does. We interpret these results in the context of the last experiment. While incorrect equivariance can cause a model to become poorly calibrated, that does not imply that correct equivariance provides a direct benefit to model calibration in the same way it does accuracy. This result is echoed in Appendix˜F, which shows similar trends for both cyclic and dihedral group order for both DESI and SDSS surveys. We note that the accuracy curves are reminiscent of Weiler & Cesa (2019); Pandya et al. (2023) with the asymptotic accuracy increase as a function of group order.
@@ 5 Invariant and Equivariant Regression Calibration
We now study calibration error in the case of invariant regression. In the process, we define a novel notion of calibration error that works for vector-valued functions. We prove that calibration error is bounded above by an expression that is analogous to the maximum $\chi^{2}$ error one can obtain averaged over all of the fibers of the uncertainty predictions. The upper bound is determined by the orbit variance of the ground truth on each fiber. We also prove that calibration error is bounded below by the minimum error over all confidence fibers in the setting where the model outputs a univariate Gaussian, which is again determined by the variance of the ground truth on each fiber.
@@ 5.1 Invariant Regression Problem Setup
Consider a function $f:X\to Y$ where $Y=\mathbb{R}^{n}$. Define a function space $\mathscr{H}$ as the set $\mathscr{H}=\{h:X\to\mathcal{M}\times\mathcal{S}\}$. Each function in the space outputs multivariate Gaussian distributions with diagonal covariance. Here, $\mathcal{M}=\mathbb{R}^{n}$ represents the space of all mean-vectors and $\mathcal{S}=\mathbb{R}_{+}^{n}$ represents the space of all variance-vectors. Denote the two outputs by $h_{\mu}$ and $h_{\sigma^{2}}$. Let $p:X\to\mathbb{R}$ be a probability density over the domain $X$. Denote the subdomain of $X$ given by the constraint $h_{\sigma^{2}}(x)=s$ as $X_{s}=\{x\in X\mid{h_{\sigma^{2}}(x)=s}\}$. Denote the fundamental domain of $G$ in $X_{s}$ as $F_{s}$. Recall smoothness hypothesis ˜2, which we assume holds for $X_{s}$ and $F_{s}$.
Next, we define a family of probability densities for each fiber of $h_{\sigma^{2}}$. Define a density over $X_{s}$ by $q_{s}:X\to\mathbb{R}$ via $q_{s}(x)=\frac{p(x_{s})}{\int_{X_{s}}p(x)dx}$. For $x\notin X_{s}$, we assume $q_{s}(x)=0$. This allows us to define the domain restricted regression error
@@
We denote the absolute value of a vector $|s|=(|s_{i}|)_{i}$ to be applied element-wise. Similarly, $\sqrt{s}$ denotes the elementwise application of the square root to $s$. Vectors have a partial ordering where $\mathbf{a}\leq\mathbf{b}$ if $a_{i}\leq b_{i}$ for all $i$. Let $D$ be the region of vectors $d$ bounded by $\mathbf{s_{1}}\leq\mathbf{d}\leq\mathbf{s_{2}}$. Define a probability density $r:\mathcal{S}\to\mathbb{R}$ such that $\mathbb{P}(h_{\sigma^{2}}(x)\in D)=\mathbb{P}(s\in D)=\int_{D}r(s)ds$. This is the push-forward of the density of $p$ over $h_{\sigma^{2}}$. We assume an analogous condition to ˜1: The input $X$ is equipped with an $|X|$ dimensional Hausdorff measure $\mathcal{H}$ so that the push-forward density $r(s)$ is defined with respect to the push-forward measure $h_{\sigma^{2}}\#(\mathcal{H})$. We note that these conditions do not need to hold for datasets with discrete input and output spaces, which we will explore in a later example.
@@
We now have the necessary machinery to define our novel generalized ENCE metric (GENCE) that applies to vector-valued functions. The goal for our learning task is that $h_{\mu}$ fits the function $f$ and $h_{\sigma^{2}}$ properly predicts confidence by minimizing GENCE (Equation˜8):
@@
We clarify why the normalization constant $\sqrt{2/\pi}$ is included in Definition˜6. If a model is well calibrated, then $\sqrt{2s/\pi}=\overline{|h_{\mu}(x)-f(x)|}$ and $s=\overline{(h_{\mu}(x)-f(x))^{2}}$ should both hold. The factor of $\sqrt{2/\pi}$ comes from Geary (1935) and is obtained by integrating the product of a normal distribution with the absolute distance between a function and the mean. In other words, if a function $h_{\sigma^{2}}$ predicts variance $s$, then we should expect the mean absolute distance between $f$ and $h_{\mu}$ to be $\sqrt{2s/\pi}$. We choose to penalizes discrepancies between $\sqrt{2s/\pi}$ and $|h_{\mu}(x)-f(x)|$ instead of $s$ and $(h_{\mu}(x)-f(x))^{2}$ because expressing the error term in terms of absolute value allows us to later apply a theorem from Wang et al. (2024).
@@
where $(\cdot)^{2}$ is applied element-wise. Unless otherwise stated, we will use the version of GENCE specified in Equation˜8. We further distinguish by using $\mathrm{GENCE}(h)$ for Equation˜8 and $\mathrm{GENCE}_{\mathrm{sq}}(h)$ for Equation˜9.
As we did for classification, we will omit the limit $\varepsilon\to 0$ in Equation˜8 and Equation˜9 for brevity, but continue to assume the expression is well defined when $r(s)\neq 0$ and $X$ has a Hausdorff measure $\mathcal{H}$.
We assume that the function space $\mathscr{H}$ is arbitrarily expressive except that it is constrained to be invariant with respect to a group $G$. That is to say, it is a universal approximator for compactly supported $G$-invariant functions.
@@ 5.2 GENCE Invariant Upper Bound
If a model $h$ is $G$-invariant, then GENCE as defined in Equation˜8 is bounded as follows
@@ 5.3 GENCE Equivariant Upper Bound
We now generalize the bounds on calibration error for invariant models (Theorem˜6) to equivariant models. While the results are similar to the invariant case, the necessary proof strategy is different. This is because elements in each fiber are no longer closed under the action of the group. That is, $G$ acting on $X_{s}\subseteq X$ may result in a vector in $X\setminus X_{s}$. Similar to Theorem˜6, our bound can be used to interpret how poorly calibrated an equivariant model can be when it is performing near its optimal capability in terms of regression error.
@@
Consider the integral in Equation˜7. Since $q_{s}(x)=0$ for $h_{\sigma^{2}}(x)\neq s$, we may write
@@
The upper bounds presented in this section are analogous to the maximum $\chi^{2}$ score one can attain averaged over all of the variances the model predicts. Consequently, fibers $\mathcal{F}_{s}$ where the regression error is small may contribute more to the upper bound than other fibers if the variance $s$ is also smaller. That being said, this $\chi^{2}$ value may not exhibit nice convergence properties.
@@
The upper bounds in Equation˜11 and Equation˜12 are not guaranteed to converge to a finite value since $\text{err}_{\text{reg}}(h,s)$ may increase as $s$ goes to infinity. It is tempting to suggest that $r(s)$ has bounded support (and therefore the integral converges). However, $r(s)$ must be supported everywhere in order for GENCE to be well defined. This suggests that bounds on GENCE are only finite when the regression error exhibits nice convergence properties, such as $\underset{s\to\infty}{\lim}\text{err}_{\text{reg}}(h,s)<\log s$. Some works (e.g., Vaicenavicius et al., 2019) circumnavigate this by defining the calibration error as the integrated density only on the set $\{s\in S:r(s)\neq 0\}$, i.e. the set where calibration error meets a necessary well definedness property.
@@
In general, it is difficult to compute the upper bound (Theorem˜6 and Theorem˜7) precisely in practice without prior knowledge on the density of the input data or the density of the push-forward. The true bound can be estimated on trained networks using sample predictions on the data to approximate $r(s)$, however, this may become intractable when the Gaussian is high dimensional. This is because the number of required samples to estimate the probability mass $r(s)$ on a small subset of $\mathbb{R}^{n}$ grows quickly with $n$. Alternatively, $r(s)$ can be analytically derived when the input distribution is well behaved and the model is simple. For example, a linear model with invertible weights will map a Gaussian distribution $p(x)$ to another Gaussian with a new mean and covariance. Despite these complications, we can consider special cases where the bound can be computed exactly.
5.4 Example of GENCE Upper Bound for Invariant Models
Example 6.
The goal of this example is to show how the GENCE bound can be computed in a small synthetic example and to show how the bound can be tight or lose depending on the norm of the variances. In this example, we consider a set $X$ of five point clouds. We calculate the GENCE upper bound assuming a minimized regression error. We assume that the function space $\mathscr{H}$ has $E(2)$-invariance. As in Section˜4.7, this dataset is adversarial to the entire function space $\mathscr{H}$ given that the ground truth function $f$ is not $E(2)$-invariant, as indicated in Figure˜5. The dataset contains some duplicates of the point clouds up to transformations in $E(2)$. Different point cloud orientations are notated with ++ and $\times$.
Now that we have established the setup, we now compute the upper bound by computing the regression error lower bound along each fiber. We start with the first fiber (rows 1 and 2 in Figure˜5). Since $h$ is $E(2)$-invariant, $h$ is able to fully fit the function $f$ on the first orbit (row 1) containing two rotated versions of $a$. Thus, the regression error is zero. For the second row, $b$ and $c$ are equally probable, and the output of $h$ that minimizes the regression error is just the average of the two lines. To compute the error on the second orbit (row 2), we also need to compute the integrated density:
$q(Gx_{s})$ $=\frac{p(b)+p(c)}{p(a(+))+p(a(\times))+p(b)+p(c)}=0.5.$
Applying Equation˜11 gives $\text{err}_{\text{reg}}(h,s_{1})=0.5\cdot\mathbb{V}_{Gx_{s_{1}}}[f]=\pi/8$. Now, as with the first orbit, the minimizing regression error on the third orbit (row 3) is zero, since there is only one element in $X$ we can fit. So, $\text{err}_{\text{reg}}(h,s_{2})=0$. Finally, that the probability of sampling an element on the fiber of $s_{1}$ is the same as the probability of sampling an element on $s_{2}$. Thus, we can appeal to Theorem˜2 and compute
[…]
We conclude that the upper bound for GENCE in our example is $1+\frac{\pi/8}{\lVert\sqrt{\frac{2}{\pi}s_{1}}\rVert_{2}^{2}}$. We can see that in the limit as $\lVert\sqrt{\frac{2}{\pi}s_{1}}\rVert_{2}^{2}$ goes to infinity, the upper bound on GENCE becomes $1$. Alternatively, in the limit as $\lVert\sqrt{\frac{2}{\pi}s_{1}}\rVert_{2}^{2}$ goes to 0, the upper bound on GENCE diverges. The interpretation of the latter is that if the model is extremely confident, then any deviations from the mean prediction represents similarly extreme miscalibration.

### v4

@@ 4.3 Improved Bounds for bi-Lipschitz Invariant Functions
Notice that the upper bounds in proposition˜1 and theorem˜4 are in terms of $r(p)$ but this density is not in general easily derivable from $q(x)$. In order to express each bound in terms of $q(x)$, we introduce extra assumptions on $h_{P}$.
From hörmander2015analysis, we have that if $h_{P}(x)$ is continuously differentiable and has gradient nowhere 0, then
@@
Assume $h_{P}(x)$ is differentiable, has gradient nowhere 0, and is $(K_{1},K_{2})$ bi-Lipschitz continuous. Let $Gx^{\diamond}$ be the orbit with the least integrated probability density, i.e., $x^{\diamond}=\underset{x\in X}{\arg\min}\int_{Gx}q(x)dx$. Then
@@
We start with the bound from theorem˜4. Substituting the expression for $r(p)$ from equation˜6 into the upper bound from theorem˜4 and using the Lipschitz constant to bound the gradient gives
@@
As in corollary˜2, the $\int_{P_{2}}dp$ term is unity in the case of binary classification. This enables us to further characterize the utility of the bound in the examples that follow.
@@ 4.4 Invariant Upper Bounds Examples
A natural question is how much tighter is the upper bound on ECE under the assumption of invariance (theorem˜4) than the bound in the unconstrained case (proposition˜1). In the case $h_{P}(x)$ is bi-Lipschitz with constants $K_{1}$ and $K_{2}$, proposition˜2 gives a relatively concrete answer; the gap is $k(Gx^{*})K_{2}\int_{P_{2}}dp\int_{Gx^{\diamond}}q(x)dx$. In order to understand the gap when the function is not assumed to be bi-Lipshitz, we consider several examples with specific $r(p)$. The examples illuminate how the tendency of a model to be uncertain can tighten the bound for functions both with and without invariance. Specifically, example˜2 considers a binary classification task where $r(p)$ is a truncated normal distribution. We consider means $\mu$ that correspond to low, medium, and high confidence.
@@
Let us now consider the unconstrained bound for three different example distributions $r(p)$, each corresponding to low, medium, and high confidence.
@@
The following example illustrates how to use the bound in corollary˜2 and how tight it is. That is, we are in the special case when we know which elements $x$ are in a given fiber $\mathcal{F}_{p}$ and the task is binary classification.
@@
Having only two confidence values may be reasonable in real world settings if there is a prevailing noise (e.g., shadow, camera artifacts, background patterns) on just one side of the field of view of a camera, leading to approximately two different confidence regions in our model output. This setup is also conceptually similar to what we will see in section˜4.7, where we will show experimentally how incorrect reflection invariance affects model calibration.
@@
Consider the same dataset as in example˜3, but assume $h$ has rotation invariance to SO(2) instead of reflection invariance. In this case, there is only one orbit and so $h_{Y}$ and $h_{P}$ each take one value. There is only one fiber $F_{p}=X$, which is the entire dataset, as illustrated in figure˜2c. The classification error is minimized when $h$ predicts each label as blue, since there are more blue labels then orange labels in our dataset. This gives an error lower bound of $0.3$, and an ECE upper bound of $0.7$. This is the same result for reflection invariance in the special case where the confidences are the same across each fiber.
Figure 2: Dataset with pointwise incorrect invariance. The first two panels indicate that the model is invariant to the action of $C_{2}$, which performs reflections over the $x$-axis. The tick mark indicates the $x=1$ point. The presence of diagonal lines in the first panel depicts that the model has $2$ confidence fibers. The last panel indicates that the model has rotation invariance. Colors represent labels in all panels.
@@ 4.6 Invariant Lower Bound Example
For Lipschitz functions $f$ and $g$ with Lipschitz constants $L_{1}$ and $L_{2}$, the composition $f(g(x))$ has Lipschitz constant $L_{2}L_{1}$ and the sum $f(x)+g(x)$ has Lipschitz constant $L_{1}+L_{2}$. Moreover, the Lipschitz constant for a linear map is given by its maximum singular value $\sigma_{\rm max}$. Finally, note that ReLU is 11-Lipschitz. Thus, the Lipschitz constant for equation˜7 is $\sigma_{\rm max}(1^{T})(\sigma_{\rm max}(I)+\sigma_{\rm max}(11^{T}))=\sigma_{\rm max}(1^{T})(1+\sigma_{\rm max}(11^{T}))$. For $n=24$, we can compute these values to be $24$ and approximately $4.9$ respectively.
@@
Thus far, we have derived a lower bound that does not depend on the size of the fibers and has no dependence on $|\nabla h_{P}|$. What is left is to calculate the integrated density on $Gx^{*}$ and the accuracy lower bound $m^{\prime}$.
Since there is only one orbit, $Gx^{*}=X$ and the integrated density over $Gx^{*}$ is 1. Moreover, $m^{\prime}$ reduces to the global accuracy lower bound. Noting that $A_{1}=\vec{a}$ will occur with probability $0.25$ gives us our accuracy lower bound $m^{\prime}$. So, we compute
@@ Results.
figure˜4 shows that ECE does not follow a roughly monotonic improvement as we increase the cyclic group order in the same way accuracy does. We interpret these results in the context of the last experiment. While incorrect equivariance can cause a model to become poorly calibrated, that does not imply that correct equivariance provides a direct benefit to model calibration in the same way it does accuracy. This result is echoed in appendix˜F, which shows similar trends for both cyclic and dihedral group order for both DESI and SDSS surveys. We note that the accuracy curves are reminiscent of Weiler and Cesa (2019); Pandya et al. (2023) with the asymptotic accuracy increase as a function of group order.
@@ 5 Invariant and Equivariant Regression Calibration
We now study calibration error in the case of invariant regression. In the process, we define a novel notion of calibration error that works for vector-valued functions. We prove that calibration error is bounded above by an expression that is analogous to the maximum $\chi^{2}$ error one can obtain averaged over all of the fibers of the uncertainty predictions. The upper bound is determined by the orbit variance of the ground truth on each fiber. We also prove that calibration error is bounded below by the minimum error over all confidence fibers in the setting where the model outputs a univariate Gaussian, which is again determined by the variance of the ground truth on each fiber. Readers interested in experimental consequences of equivariance on uncertainty in the regression setting may skip to section˜6.
@@ 5.1 Invariant Regression Problem Setup
Consider a function $f:X\to Y$ where $Y=\mathbb{R}^{n}$. Define a function space $\mathscr{H}$ as the set $\mathscr{H}=\{h\colon X\to\mathcal{M}\times\mathcal{S}\}$. Each function in the space outputs multivariate Gaussian distributions with diagonal covariance. Here, $\mathcal{M}=\mathbb{R}^{n}$ represents the space of all mean-vectors and $\mathcal{S}=\mathbb{R}_{+}^{n}$ represents the space of all variance-vectors. Denote the two outputs by $h_{\mu}$ and $h_{\sigma^{2}}$ and let $p\colon X\to\mathbb{R}$ be a probability density over the domain $X$. Denote the subdomain of $X$ given by the constraint $h_{\sigma^{2}}(x)=s$ as $X_{s}=\{x\in X\mid{h_{\sigma^{2}}(x)=s}\}$. Denote the fundamental domain of $G$ in $X_{s}$ as $F_{s}$. Recall smoothness and separability hypothesis in ˜2, which we assume holds for $X_{s}$ and $F_{s}$.
Next, we define a family of probability densities for each fiber of $h_{\sigma^{2}}$. Define a density over $X_{s}$ by $q_{s}\colon X\to\mathbb{R}$ via $q_{s}(x)=\frac{p(x_{s})}{\int_{X_{s}}p(x)dx}$. For $x\notin X_{s}$, we assume $q_{s}(x)=0$. This allows us to define the domain restricted regression error
@@
We denote the absolute value of a vector $|s|=(|s_{i}|)_{i}$ to be applied element-wise. Similarly, $\sqrt{s}$ denotes the elementwise application of the square root to $s$. Vectors have a partial ordering where $\mathbf{a}\leq\mathbf{b}$ if $a_{i}\leq b_{i}$ for all $i$. Let $D$ be the region of vectors $d$ bounded by $\mathbf{s_{1}}\leq\mathbf{d}\leq\mathbf{s_{2}}$. Define a probability density $r\colon\mathcal{S}\to\mathbb{R}$ such that $\mathbb{P}(h_{\sigma^{2}}(x)\in D)=\mathbb{P}(s\in D)=\int_{D}r(s)ds$. This is the push-forward of the density of $p$ over $h_{\sigma^{2}}$. We assume an analogous condition to ˜1: The input $X$ is equipped with an $|X|$ dimensional Hausdorff measure $\mathcal{H}$ so that the push-forward density $r(s)$ is defined with respect to the push-forward measure $h_{\sigma^{2}}\#(\mathcal{H})$. We note that these conditions do not need to hold for datasets with discrete input and output spaces, which we will explore in a later example.
@@
We now have the necessary machinery to define our novel Generalized ENCE (GENCE) metric that applies to vector-valued functions. The goal for our learning task is that $h_{\mu}$ fits the function $f$ and $h_{\sigma^{2}}$ properly predicts confidence by minimizing GENCE (equation˜9):
@@
We clarify why the normalization constant $\sqrt{2/\pi}$ is included in definition˜6. If a model is well calibrated, then $\sqrt{2s/\pi}=\overline{|h_{\mu}(x)-f(x)|}$ and $s=\overline{(h_{\mu}(x)-f(x))^{2}}$ should both hold. The factor of $\sqrt{2/\pi}$ comes from Geary (1935) and is obtained by integrating the product of a normal distribution with the absolute distance between a function and the mean. In other words, if a function $h_{\sigma^{2}}$ predicts variance $s$, then we should expect the mean absolute distance between $f$ and $h_{\mu}$ to be $\sqrt{2s/\pi}$. We choose to penalize discrepancies between $\sqrt{2s/\pi}$ and $|h_{\mu}(x)-f(x)|$ instead of $s$ and $(h_{\mu}(x)-f(x))^{2}$ because expressing the error term in terms of absolute value allows us to later apply a theorem from Wang et al. (2024).
@@
where $(\cdot)^{2}$ is applied element-wise. Unless otherwise stated, we will use the version of GENCE specified in equation˜9. We further distinguish by using $\mathrm{GENCE}(h)$ for equation˜9 and $\mathrm{GENCE}_{\mathrm{sq}}(h)$ for equation˜10.
As we did for classification, we will omit the limit $\varepsilon\to 0$ in equation˜9 and equation˜10 for brevity, but continue to assume the expression is well defined when $r(s)\neq 0$ and $X$ has a Hausdorff measure $\mathcal{H}$. We assume that the function space $\mathscr{H}$ is arbitrarily expressive except that it is constrained to be invariant with respect to a group $G$. That is to say, it is a universal approximator for compactly supported $G$-invariant functions.
We adopt $\mathrm{GENCE}(h)$ and $\mathrm{GENCE}_{\mathrm{sq}}(h)$ as calibration metrics because it enables us to continue our strategy of bounding calibration errors by noting that the calibration function is equivalent to a fiber-wise error, but note that alternative frameworks such as coverage based metrics are also compelling. In particular, extending Dobriban and Yu (2025) to cases of symmetry mismatch would serve as a complement to this work.
@@ 5.2 GENCE Invariant Upper Bound
If a model $h$ is $G$-invariant, then GENCE as defined in equation˜9 is bounded as follows
@@ 5.3 GENCE Equivariant Upper Bound
We now generalize the bounds on calibration error for invariant models (theorem˜6) to equivariant models. While the results are similar to the invariant case, the proof strategy is different. This is because the fibers are no longer closed under the action of the group. That is, $G$ acting on $X_{s}\subseteq X$ may result in a vector in $X\setminus X_{s}$. Similar to theorem˜6, our bound can be used to interpret how poorly calibrated an equivariant model can be when it is performing near its optimal capability in terms of regression error.
@@
Consider the integral in equation˜8. Since $q_{s}(x)=0$ for $h_{\sigma^{2}}(x)\neq s$, we may write
@@
The upper bounds presented in this section are analogous to the maximum $\chi^{2}$ score one can attain averaged over all of the variances the model predicts. Consequently, fibers $\mathcal{F}_{s}$ where the regression error is small may contribute more to the upper bound than other fibers if the variance $s$ is also smaller. That being said, this $\chi^{2}$ value may not exhibit favorable convergence properties.
@@
The upper bounds in equation˜12 and Equation˜13 are not guaranteed to converge to a finite value since $\text{err}_{\text{reg}}(h,s)$ may increase as $s$ goes to infinity. It is tempting to suggest that $r(s)$ has bounded support (and therefore the integral converges). However, $r(s)$ must be supported everywhere in order for GENCE to be well defined. This suggests that bounds on GENCE are only finite when the regression error exhibits nice convergence properties, such as $\underset{s\to\infty}{\lim}\text{err}_{\text{reg}}(h,s)<\log s$. Some works (e.g., Vaicenavicius et al., 2019) circumnavigate this by defining the calibration error as the integrated density only on the set $\{s\in S\colon r(s)\neq 0\}$, i.e., the set where calibration error meets a necessary well-definedness property.
@@
In practice, it is difficult to compute the upper bound (theorem˜6 and theorem˜7) precisely without prior knowledge on the density of the input data or the density of the push-forward. The true bound can be estimated for trained networks using sample predictions on the data to approximate $r(s)$, however, this may become intractable when the Gaussian is high dimensional. This is because the number of required samples to estimate the probability mass $r(s)$ on a small subset of $\mathbb{R}^{n}$ grows quickly with $n$. Alternatively, $r(s)$ can be analytically derived when the input distribution is well behaved and the model is simple. For example, a linear model with invertible weights will map a Gaussian distribution $p(x)$ to another Gaussian with a new mean and covariance. Despite these complications, we can consider special cases where the bound can be computed exactly.